OpenStudy (anonymous):

An object will take t seconds to drop d ft, where t = square root d/16 . What distance will the object drop in square root 19/2 sec

OpenStudy (anonymous):

\[t=\sqrt{d/16}\] We know the time, so plug it in for t and solve for d: \[\sqrt{19/2} = \sqrt{d/16}\] \[19/2 = d/16\] \[d= (19)(16)/2 = 19(8)=152 ft\]

OpenStudy (anonymous):

You made it looks easy! Thank you!

OpenStudy (anonymous):

You're welcome, glad I could help.